OpenStudy (anonymous):

solve 5/4x+1/8x=5/8+x the solution is x =

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

x=5/3

OpenStudy (anonymous):

(10x + x)/8 = (5+8x)/8 11x -8x = 5 3x = 5 -> x =5/3